I need to find out how to solve a two variables equations cause I am having a very hard time trying to figure out what to do with both variables.
a+18=7a

3w+6=4w

Found 2 solutions by rothauserc, rfer:
Answer by rothauserc(4718)

1) a + 18 = 7a
subtract a from both sides of =
18 = 6a
divide both sides of = by 6
3 = a
****************************************
2) 3w + 6 = 4w
subtract 3w from both sides of =
6 = w
***************************************
check answers
1) 3 + 18 = 7*3
21 = 21
answer checks
***************************************
2) 3*6 + 6 = 4*6
24 = 24
answer checks

Answer by rfer(16322)

a+18=7a
a-7a=-18
-6a=-18
a=-18/-6
a=+3
----------------
3w+6=4w
3w-4w=-6
-w=-6
w=+6